# A Fractal-based Complex Belief Entropy for Uncertainty Measure in Complex Evidence Theory

### v1

@@ Abstract
Complex evidence theory, as a generalized D-S evidence theory, has attracted academic attention because it can well express uncertainty by means of complex basic belief assignment (CBBA), and realize uncertainty reasoning by complex combination rule. However, the uncertainty measurement in complex evidence theory is still an open issue. In order to make better decisions, a complex pignistic belief transformation (CPBT) method has been proposed to assign CBBAs of multi-element focal elements to subsets. The essence of CPBT is the redistribution of complex mass function by means of the concept of fractal. In this paper, based on fractal theory, experimental simulation and analysis have been carried out on the generation process of CPBT in time dimension. Then, a new fractal-based complex belief (FCB) entropy is proposed to measure the uncertainty of CBBA. Finally, the properties of FCB entropy are analyzed, and several examples are used to verify its effectiveness.
@@ I Introduction
Fractal widely exists in nature and was first proposed to measure the length of the coast. The fractal theory established today is a very popular and active new theory and discipline. With people’s attention to fractal theory, it has been widely used in many different fields. The most basic feature of fractal theory is to study objective things with fractal dimension, which provides a new perspective for the analysis of complex information systems [1]. In addition, the self-similarity of fractal also has the prospect of application in information analysis.
Because of the need for accurate decision-making, uncertainty measurement is an essential issue in information analysis theory. In order to solve this problem, many relevant theories have been proposed, such as Z-number [2], intuitionistic fuzzy set [3, 4], rough set [5, 6], random permutation set [7], Dempster-Shafer evidence theory (DSET), etc [8, 9, 10]. Owing to its advantages in probability distribution of uncertain information, DSET and its branches [11] are widely used in many fields, such as decision-making [12, 13, 14, 15, 16, 17], pattern recognition [18, 19], multi-source information fusion [20, 21, 22], classification decision [23, 24], reasoning [25, 26, 27], fault diagnosis [28], etc [29, 30, 31, 32, 33, 34].
DSET is a generalization of probability theory. It assigns the probability to the power set of the element to represent the corresponding information, and uses the mass function to represent the probability of the power set, namely basic belief assignment (BBA). How to understand and measure uncertainty has always been a crucial issue in DSET.
BBA is not only affected by the probability of the event itself, but also by the uncertainty of the set allocation. Therefore, the uncertainty of DEST usually consists of two parts: discord and non-specificity. Discord represents the conflict between different elements in the framework, and non-specificity represents the uncertainty generated during the distribution of BBA. At present, there are two main types of BBA uncertainty. Firstly, based on Shannon entropy and other ideas, some methods use entropy [35] to measure uncertainty, such as Hohle entropy [höhle1982entropy], weighted Hartley entropy [36], Pal et al.’s entropy [37], Jousselme et al.’s entropy [38], the entropy of Jiroušek and Shenoy [39], Pan et al.’s entropy [40], Deng’s entropy [41], etc [42, 43]. Comparatively, in order to avoid the difference between DSET and probability theory due to the discernment of framework, the non-entropy method uses belief function and likelihood function to directly define on the framework, such as the uncertainty measure of Song and Wang [44], the distance-based total uncertainty measures of Yang and Han [45, 46], the total uncertainty measure of interval-valued belief structures [47], etc [48, 49, 50].
However, the mass function of DSET is defined in the real number field, which leads to limitations when considering the fluctuations of data at a specific phase. DSET cannot handle the uncertain information in the complex field well. Therefore, the methods mentioned above cannot describe the information in the complex field well. To solve this issue, the complex evidence theory (CET), also known as generalized DSET, was proposed by Xiao [51, 52]. Different from DSET, complex basic belief assignment (CBBA) in CET can well express information in the complex field, and complex evidence combination rules can be well applied to the uncertainty reasoning based on the complex numbers [53, 54, 55]. Specifically, CET has shown great potential and advantages for information modeling in the quantum field [56, 57]. In short, CET provides a new perspective for uncertain information reasoning, and one of the most noteworthy issues in CET is the uncertainty measure of CBBA.
Whether DSET or CET, the ultimate goal is to obtain the probability of mutually exclusive single elements in the framework. In classical DSET, several probability transformation methods, such as pignisitic probability transformation (PPT) [58], plausibility transformation method (PTM) [59] and other probability transformation methods [34] are proposed. In CET, to build a connection between CBBA and probability, a complex pignistic belief transformation (CPBT) [56] is proposed to convert complex mass assignments in CET to probability distribution. However, these methods only give the final results, and do not propose specific production process. Therefore, the evaluation of these methods is not comprehensive.
Recently, the research on the relationship of fractal theory and evidence theory has attracted academic attention. Inspired by [60], a CPBT generation process based on fractal idea is proposed, and on the basis of this process, a fractal-based complex belief (FCB) entropy is proposed to measure the total uncertainty of CBBA. Then, the feasibility of FCB entropy is verified by theoretical analysis and mathematical proof, and the performance of FCB entropy is demonstrated by concrete examples.
@@
The remainder of the paper is structured as follows. In Section II, some basic concepts of DSET and CET are briefly reviewed, and then fractal theory and some common entropies are briefly introduced. In Section III, a CPBT generation process based on fractal idea is proposed firstly, and then FCB entropy is proposed based on the above process. Then the properties of FCB entropy are analyzed and proved in Section IV. Then several examples are used to verify the effectiveness of FCB entropy in Section V. Section VI is focused on conclusions and future prospects.
@@ II Preliminaries
In Section 2, the basic theories of DSET and CET are briefly reviewed, and then fractal theory and some common entropies in evidence theory are introduced.
@@ II-A Dempster-Shafer evidence theory
Definition II.1.
(Frame of Discernment). Let a set be exhaustible and the elements in it are mutually exclusive, which can be denoted as $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$. Then the set $\Theta$ is called a frame of discernment (FoD). The power set of $\Theta$ is denoted as ${{2}^{\Theta}}=\{\varnothing,\{{{e}_{1}}\},\{{{e}_{2}}\},\cdots\{{{e}_{1}},{{e}_{2}}\},\cdots,\Theta\}$, where $\varnothing$ is an empty set.
Definition II.2.
(Mass Function). Let $m$ be a mss fuction in FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, which is a mapping:$m:{{2}^{\Theta}}\to\mathbb{R}$ satisfying
$\small m(\varnothing)=0,$ (1)
$\small\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})}=1.$ (2)
@@
Definition II.3.
(Pignistic Probability Transformation). Given a FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$ composed of n elements and the corresponding BBA $m$, its pignistic probability transformation (PPT) is defined as:
$\small BetP({{e}_{i}})=\sum\limits_{{{e}_{i}}\in{{A}_{i}}and{{A}_{i}}\in{{2}^{\Theta}}}{\frac{m({{A}_{i}})}{\left|{{A}_{i}}\right|}},$ (3)
where $\left|{{A}_{i}}\right|$ represents the cardinality of ${{A}_{i}}$.
@@ II-B Complex evidence theory
Definition II.4.
(Complex Mass Function). Let $\mathbb{M}$ be a complex mass function (CMF) in FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, which is a mapping: $\mathbb{M}:{{2}^{\Theta}}\to\mathbb{C}$ satisfying
$\small\mathbb{M}(\varnothing)=0,$ (4)
$\small\mathbb{M}({{A}_{i}})=\mathbf{m}({{A}_{i}}){{e}^{i\theta({{A}_{i}})}},{{A}_{i}}\in{{2}^{\Theta}},$ (5)
@@
$\small\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{\mathbb{M}({{A}_{i}})}=1,$ (6)
@@
$\small\mathbb{M}({{A}_{i}})=x+yi,{{A}_{i}}\in{{2}^{\Theta}},$ (7)
$\small\left|\mathbb{M}({{A}_{i}})\right|=\mathbf{m}({{A}_{i}})=\sqrt{{{x}^{2}}+{{y}^{2}}}.$ (8)
@@
Definition II.5.
(Commitment Degree). The commitment degree $\mathbb{C}om({{A}_{i}})$ represent the support degree for ${{A}_{i}}$, which is expressed as
$\small\mathbb{C}om({{A}_{i}})=\frac{\left|\mathbb{M}({{A}_{i}})\right|}{\sum\nolimits_{{{B}_{i}}\in{{2}^{\Theta}}}{\left|\mathbb{M}({{B}_{i}})\right|}}.$ (9)
Definition II.6.
@@
$\small CBet({{A}_{i}})=\sum\limits_{{{A}_{i}}\subseteq{{B}_{i}}\in{{2}^{\Theta}}}{\frac{\mathbb{M}({{B}_{i}})}{\left|{{B}_{i}}\right|}},$ (10)
@@
Definition II.7.
(Interference Effect). Interference effect refers to the residual term generated when calculating the added complex modulus, which is defined as
@@
Definition II.8.
(CBBA Exponential Negation). Given a FoD $\Theta=\{{{e}_{1}}$, ${{e}_{2}}$, $\cdots$, ${{e}_{n}}\}$ and the power set of $\Theta$ is ${{2}^{\Theta}}=\{\varnothing,\{{{e}_{1}}\},\{{{e}_{2}}\},\cdots,\{{{e}_{1}},{{e}_{2}}\},\cdots,\Theta\}$. Let a CBBA $\mathbb{M}$ be defined on ${{2}^{\Theta}}$. Then the CBBA exponential negation is defined by
$\small\overline{\mathbb{M}}({{A}_{i}})=\frac{\sum\limits_{B\neq\Theta,{{B}_{i}}\in{{2}^{\Theta}}}{{{e}^{-\mathbb{M}({{B}_{i}})\cdot\left|{{A}_{i}}\cap\overline{{{B}_{i}}}\right|}}}}{\sum\limits_{{{A}_{i}}\neq\varnothing,{{A}_{i}}\in{{2}^{\Theta}}}{\sum\limits_{{{B}_{i}}\neq\Theta,{{B}_{i}}\in{{2}^{\Theta}}}{{{e}^{-\mathbb{M}({{B}_{i}})\cdot\left|{{A}_{i}}\cap\overline{{{B}_{i}}}\right|}}}}},A\neq\varnothing,$ (12)
@@ II-C Entropy
Definition II.9.
(Shannon Entropy). Given the probability distribution of all events in a framework $P=\{{{p}_{1}},\cdots,{{p}_{n}}\}$, then the Shannon entropy is defined as below:
$\small{{E}_{s}}(P)=-\sum\limits_{i=1}^{n}{{{p}_{i}}\log{{p}_{i}}}.$ (13)
Definition II.10.
(Weighted Hartley Entropy). Weighted Hartley entropy is extended from Hartley information formula, which is defined as follows:
$\small{{E}_{H}}(m)=\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})\log\left|{{A}_{i}}\right|}.$ (14)
Definition II.11.
(Pal et al.’s Entropy). Given a mass function $m$ in an $n$-element FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, then the Pal et al.’s entropy is defined as below:
$\small{{E}_{p}}(m)=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})\log(\frac{m({{A}_{i}})}{\left|{{A}_{i}}\right|})}.$ (15)
Definition II.12.
(Deng Entropy). Given a mass function $m$ in an $n$-element FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, then the Deng entropy is defined as below:
$\small{{E}_{d}}(m)=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})\log\frac{m({{A}_{i}})}{{{2}^{\left|{{A}_{i}}\right|}}-1}},$ (16)
Definition II.13.
(Zhou et al.s’ Entropy). Given a mass function $m$ in an $n$-element FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, then the Zhou et al.s’ entropy is defined as below:
$\small{{E}_{z}}(m)=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})\log(\frac{m({{A}_{i}})}{{{2}^{\left|A\right|}}-1}{{e}^{\frac{\left|A\right|-1}{\left|X\right|}}})}.$ (17)
Definition II.14.
(Cui et al.s’ Entropy). Given a mass function $m$ in an $n$-element FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, then the Cui et al.s’ entropy [62] is defined as below:
$\small{{E}_{c}}(m)=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})\log}\left(\frac{m({{A}_{i}})}{{{2}^{\left|{{A}_{i}}\right|}}-1}{{e}^{\sum\limits_{\begin{smallmatrix}{{B}_{i}}\in{{2}^{\Theta}}\\ {{B}_{i}}\neq{{A}_{i}}\end{smallmatrix}}{\frac{\left|{{A}_{i}}\cap{{B}_{i}}\right|}{{{2}^{\left|\Theta\right|}}-1}}}}\right).$ (18)
Definition II.15.
(FB Entropy). Given a mass function $m$ in an $n$-element FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, then the FB entropy is defined as below:
$\small{{E}_{FB}}(m)=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{{{m}_{F}}({{A}_{i}})\log{{m}_{F}}({{A}_{i}})},$ (19)
@@
$\small{{m}_{F}}({{A}_{i}})=\frac{m({{A}_{i}})}{{{2}^{\left|{{A}_{i}}\right|}}-1}+\sum\limits_{{{A}_{i}}\subset{{A}_{j}}}{\frac{m({{A}_{j}})}{{{2}^{\left|{{A}_{j}}\right|}}-1}},$ (20)

### v2

@@ Abstract
Complex Evidence Theory (CET), an extension of the traditional D-S evidence theory, has garnered academic interest for its capacity to articulate uncertainty through Complex Basic Belief Assignment (CBBA) and to perform uncertainty reasoning using complex combination rules. Nonetheless, quantifying uncertainty within CET remains a subject of ongoing research. To enhance decision-making, a method for Complex Pignistic Belief Transformation (CPBT) has been introduced, which allocates CBBAs of multi-element focal elements to subsets. CPBT’s core lies in the fractal-inspired redistribution of the complex mass function. This paper presents an experimental simulation and analysis of CPBT’s generation process along the temporal dimension, rooted in fractal theory. Subsequently, a novel Fractal-Based Complex Belief (FCB) entropy is proposed to gauge the uncertainty of CBBA. The properties of FCB entropy are examined, and its efficacy is demonstrated through various numerical examples and practical application.
@@ I Introduction
Precise decision-making necessitates the quantification of uncertainty, dynamic and reliability analysis, a critical aspect within information analysis theory [1, 2, 3, 4, 5, 6, 7]. To address this, various theories have been introduced, including the extended Z-number [8], D number [9], rough sets [10, 11], random permutation sets [12], Dempster-Shafer evidence theory (DSET) [13, 14] and so on [15]. DSET, in particular, and its extensions [16, 17] are favored for their nuanced handling of uncertain information probabilities and have been extensively applied across numerous domains. These applications span decision-making [18, 19], multi-source information fusion [20], classification [21, 22, 23], reasoning [24, 25], and diagnosis [26, 27], among others [28, 29, 30].
DSET serves as an extension of traditional probability theory, wherein probabilities are assigned to the power set of elements to encapsulate the represented information [31, 32]. This is achieved through the use of a mass function, which quantifies the probability of each power set, known as the basic belief assignment (BBA). Grappling with the concepts of understanding and quantifying uncertainty lies at the heart of DSET [33, 34]. BBA in DSET is influenced not only by the inherent probability of an event but also by the uncertainty associated with the allocation of sets. However, the mass function in DSET operates within the real number field, presenting limitations when capturing the nuances of data fluctuations at specific stages. DSET falls short in managing uncertain information in the complex number field [35]. Consequently, the aforementioned methods struggle to accurately represent information within this domain. To address this, Xiao introduced Complex Evidence Theory (CET), also known as Generalized DSET [36, 37]. Unlike DSET, CET’s Complex Basic Belief Assignment (CBBA) effectively encapsulates information in the complex field, and its complex evidence combination rules are adept at handling uncertainty reasoning with complex numbers [38, 39]. Notably, CET has demonstrated significant potential and advantages for information modeling in quantum fields [40, 41, 42, 43, 44]. In essence, CET offers a novel framework for uncertain information reasoning, with the uncertainty measure of CBBA being one of its most critical aspects.
To be specific, the uncertainty in DSET is typically composed of discord and non-specificity. Discord reflects the conflicts among different elements within the framework [45], while non-specificity captures the uncertainty that arises during the distribution of BBA [46]. Currently, two primary approaches are used to quantify BBA uncertainty. The first approach leverages entropy-based methods, drawing on concepts such as Shannon entropy. This category includes various entropy measures like Hohle entropy [47], weighted Hartley entropy [48], Pal et al.’s entropy [49], Jousselme et al.’s entropy [50], Deng’s entropy [51], Jiroušek and Shenoy’s entropy [52], Cui and Deng’s Plausibility entropy [53], among others [54, 55]. In contrast, the non-entropy approach sidesteps the potential divergence between DSET and probability theory by directly defining uncertainty measures on the framework using belief and likelihood functions. Examples include Song and Wang’s uncertainty measure [56], Yang and Han’s distance-based total uncertainty measures [57, 58], and the total uncertainty measure for interval-valued belief structures [59].
While the DSET and its extensions have been instrumental in advancing the field of uncertainty quantification, the existing methods for uncertainty measurement within CET are still in their nascent stages. Most of the research efforts in CET have been building upon the foundational work established by classical DSET. This includes the exploration of the uncertainty measures that are compatible with the complex nature of CBBAs, which is a departure from the traditional real-number-based BBAs in DSET. The existing complex-valued Deng entropy [60] transforms Deng entropy into complex form, and quantum belief entropy [43] uses density matrix to measure the uncertainty of CBBA. Although these methods use complex numbers for uncertainty modeling, they do not fully utilize the influence of the intersection of different focal elements and maximize the use of complex number representation information.
To address limitations in uncertainty measurements like complex-valued Deng entropy and quantum belief entropy within CET, fractal theory offers a compelling approach. Fractals, known for their self-similarity and ability to analyze complex systems, have found applications across various disciplines. The fractal dimension provides new insights into complex information systems, aligning with CET’s need for advanced uncertainty quantification. Drawing inspiration from [61], a novel fractal-based Complex Pignistic Belief Transformation (CPBT) process has been introduced, leading to the development of Fractal-Based Complex Belief (FCB) entropy. This new metric, FCB entropy, is designed to capture the overall uncertainty in CBBAs more effectively. Theoretical analysis and practical applications substantiate the viability of FCB entropy, showcasing its potential to enhance decision-making and system understanding within CET by leveraging the full capacity of complex numbers.
@@
The remainder of the paper is structured as follows. In Section II, some basic concepts of DSET and CET are briefly reviewed, and then fractal theory and some common entropies are briefly introduced. In Section III, a CPBT generation process based on fractal idea is proposed firstly, and then FCB entropy is proposed based on the above process. Then the properties of FCB entropy are analyzed and proved in Section IV. Then numerical example and a algorithm with its practical application are used to verify the effectiveness of FCB entropy in Section V. Section VI is focused on conclusions and future prospects.
@@ II Preliminaries
How to model and measure the uncertainty and dynamics of the system [62, 63, 64, 65] have attracted a lot of attentions. In this section, the basic theories of DSET and CET are briefly reviewed, and then fractal theory and some common entropies in evidence theory are introduced.
@@ II-A Dempster-Shafer evidence theory
Definition 1.
(Frame of Discernment). Let a set be exhaustible and the elements in it are mutually exclusive, which can be denoted as $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$. Then the set $\Theta$ is called a frame of discernment (FoD). The power set of $\Theta$ is denoted as ${{2}^{\Theta}}=\{\varnothing,\{{{e}_{1}}\},\{{{e}_{2}}\},\cdots\{{{e}_{1}},{{e}_{2}}\},\cdots,\Theta\}$, where $\varnothing$ is an empty set [13, 14].
Definition 2.
(Mass Function). Let $m$ be a mss fuction in FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, which is a mapping:$m:{{2}^{\Theta}}\to\mathbb{R}$ [13, 14] satisfying
$m(\varnothing)=0,$ (1)
$\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})}=1.$ (2)
@@ II-B Complex evidence theory
Definition 3.
(Complex Mass Function). Let $\mathbb{M}$ be a complex mass function (CMF) in FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, which is a mapping: $\mathbb{M}:{{2}^{\Theta}}\to\mathbb{C}$ satisfying [36, 37]
$\mathbb{M}(\varnothing)=0,$ (3)
$\mathbb{M}({{A}_{i}})=\mathbf{m}({{A}_{i}}){{e}^{i\theta({{A}_{i}})}},{{A}_{i}}\in{{2}^{\Theta}},$ (4)
@@
$\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{\mathbb{M}({{A}_{i}})}=1,$ (5)
@@
$\mathbb{M}({{A}_{i}})=x+yi,{{A}_{i}}\in{{2}^{\Theta}},$ (6)
$\left|\mathbb{M}({{A}_{i}})\right|=\mathbf{m}({{A}_{i}})=\sqrt{{{x}^{2}}+{{y}^{2}}}.$ (7)
@@
Definition 4.
(Commitment Degree). The commitment degree $\mathbb{C}om({{A}_{i}})$ represents the support degree for ${{A}_{i}}$, which is expressed as [36, 37]
$\mathbb{C}om({{A}_{i}})=\frac{\left|\mathbb{M}({{A}_{i}})\right|}{\sum\nolimits_{{{B}_{i}}\in{{2}^{\Theta}}}{\left|\mathbb{M}({{B}_{i}})\right|}}.$ (8)
Definition 5.
@@
$CBet({{A}_{i}})=\sum\limits_{{{A}_{i}}\subseteq{{B}_{i}}\in{{2}^{\Theta}}}{\frac{\mathbb{M}({{B}_{i}})}{\left|{{B}_{i}}\right|}},$ (9)
@@
Definition 6.
(Interference Effect). Interference effect refers to the residual term generated when calculating the added complex modulus, which is defined as [66]
@@
Definition 7.
(CBBA Exponential Negation). Given a FoD $\Theta=\{{{e}_{1}}$, ${{e}_{2}}$, $\cdots$, ${{e}_{n}}\}$ and the power set of $\Theta$ is ${{2}^{\Theta}}=\{\varnothing,\{{{e}_{1}}\},\{{{e}_{2}}\},\cdots,\{{{e}_{1}},{{e}_{2}}\},\cdots,\Theta\}$. Let a CBBA $\mathbb{M}$ be defined on ${{2}^{\Theta}}$. Then the CBBA exponential negation is defined by [39]
$\overline{\mathbb{M}}({{A}_{i}})=\frac{\sum\limits_{B\neq\Theta,{{B}_{i}}\in{{2}^{\Theta}}}{{{e}^{-\mathbb{M}({{B}_{i}})\cdot\left|{{A}_{i}}\cap\overline{{{B}_{i}}}\right|}}}}{\sum\limits_{{{A}_{i}}\neq\varnothing,{{A}_{i}}\in{{2}^{\Theta}}}{\sum\limits_{{{B}_{i}}\neq\Theta,{{B}_{i}}\in{{2}^{\Theta}}}{{{e}^{-\mathbb{M}({{B}_{i}})\cdot\left|{{A}_{i}}\cap\overline{{{B}_{i}}}\right|}}}}},A\neq\varnothing,$ (11)
@@
Definition 8.
Let ${\mathbb{M}_{1}}$ and ${\mathbb{M}_{2}}$ are two CBBAs defined in the FOD $\Omega$. The complex evidence combination rule is defined as [36, 37]
$\mathbb{M}(C)=\left\{\begin{aligned} &\frac{1}{1-K}\sum\limits_{A\cap B=C}{{\mathbb{M}_{1}}(A){\mathbb{M}_{2}}(B)},&C\neq\varnothing,\\ &0,&C=\varnothing,\\ \end{aligned}\right.$ (12)
with
$\mathbb{K}=\sum\limits_{A\cap B=\varnothing}{{\mathbb{M}_{1}}(A){\mathbb{M}_{2}}(B)},$ (13)
where $\mathbb{K}$ is the conflict coefficient between CBBAs ${\mathbb{M}_{1}}$ and ${\mathbb{M}_{2}}$.
@@ II-C Entropy
Definition 9.
(Shannon Entropy). Given the probability distribution of all events in a framework $P=\{{{p}_{1}},\cdots,{{p}_{n}}\}$, then the Shannon entropy is defined as below [67]:
${{E}_{s}}(P)=-\sum\limits_{i=1}^{n}{{{p}_{i}}\log{{p}_{i}}}.$ (14)
Definition 10.
(Weighted Hartley Entropy). Weighted Hartley entropy is extended from Hartley information formula, which is defined as follows [68]:
${{E}_{H}}(m)=\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})\log\left|{{A}_{i}}\right|}.$ (15)
Definition 11.
(Pal et al.’s Entropy). Given a mass function $m$ in an $n$-element FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, then the Pal et al.’s entropy is defined as below [49]:
${{E}_{p}}(m)=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})\log(\frac{m({{A}_{i}})}{\left|{{A}_{i}}\right|})}.$ (16)
Definition 12.
(Deng Entropy). Given a mass function $m$ in an $n$-element FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, then the Deng entropy is defined as below [51]:
${{E}_{d}}(m)=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})\log\frac{m({{A}_{i}})}{{{2}^{\left|{{A}_{i}}\right|}}-1}},$ (17)
Definition 13.
(Zhou et al.s’ Entropy). Given a mass function $m$ in an $n$-element FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, then the Zhou et al.s’ entropy is defined as below [69]:
${{E}_{z}}(m)=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})\log(\frac{m({{A}_{i}})}{{{2}^{\left|A\right|}}-1}{{e}^{\frac{\left|A\right|-1}{\left|X\right|}}})}.$ (18)
Definition 14.
(Cui et al.s’ Entropy). Given a mass function $m$ in an $n$-element FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, then the Cui et al.s’ entropy is defined as below [70]:
${{E}_{c}}(m)=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})\log}\left(\frac{m({{A}_{i}})}{{{2}^{\left|{{A}_{i}}\right|}}-1}{{e}^{\sum\limits_{\begin{smallmatrix}{{B}_{i}}\in{{2}^{\Theta}}\\ {{B}_{i}}\neq{{A}_{i}}\end{smallmatrix}}{\frac{\left|{{A}_{i}}\cap{{B}_{i}}\right|}{{{2}^{\left|\Theta\right|}}-1}}}}\right).$ (19)
Definition 15.
(FB Entropy). Given a mass function $m$ in an $n$-element FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$, then the FB entropy is defined as below [61]:
${{E}_{FB}}(m)=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{{{m}_{F}}({{A}_{i}})\log{{m}_{F}}({{A}_{i}})},$ (20)
@@
${{m}_{F}}({{A}_{i}})=\frac{m({{A}_{i}})}{{{2}^{\left|{{A}_{i}}\right|}}-1}+\sum\limits_{{{A}_{i}}\subset{{A}_{j}}}{\frac{m({{A}_{j}})}{{{2}^{\left|{{A}_{j}}\right|}}-1}},$ (21)
Definition 16.
(Generalized entropy with its maximum). Because the entropies mentioned above are defined in the real number field, while CBBA is defined in the complex number field. The above entropies cannot be directly used in CET. A generalized entropy in CET is defined as follows [39]:
$\mathbb{E}(\mathbb{M})=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{\mathbb{C}om({{A}_{i}})\cdot\log(\frac{\mathbb{C}om({{A}_{i}})}{h({{A}_{i}})})}.$ (22)
when the generalized entropy is used in BBA, it can be expressed as
$E(m)=-\sum\limits_{{{A}_{i}}\in{{2}^{\Theta}}}{m({{A}_{i}})\cdot\log(\frac{m({{A}_{i}})}{h({{A}_{i}})})}.$ (23)
Definition 17.
(Complex-valued Deng entropy). Given a CBBA $\mathbb{CM}$ defined on a FOD $\mathbb{E}$, then complex-valued Deng entropy is defined as follows [60]:
$C{{E}_{d}}$ $=\left\|-\sum\limits_{A\subseteq\mathbb{E}}{\left\|\mathbb{CM}\left(A\right)\right\|\ln\frac{\mathbb{CM}\left(A\right)}{{{2}^{\left|A\right|}}-1}}\right\|$ (24)
$=\left\|-\sum\limits_{A\subseteq\mathbb{E}}{\left\|m\left(A\right){{e}^{i\pi{{\theta}_{A}}}}\right\|\ln\frac{m\left(A\right){{e}^{i\pi{{\theta}_{A}}}}}{{{2}^{\left|A\right|}}-1}}\right\|$
$=\left\|-\sum\limits_{A\subseteq\mathbb{E}}{m\left(A\right)\ln\frac{m\left(A\right){{e}^{i\pi{{\theta}_{A}}}}}{{{2}^{\left|A\right|}}-1}}\right\|$
Definition 18.
(Quantum belief entropy) Let a CBBA $\mathbb{M}$ define on a FOD $\Theta=\left\{{{\theta}_{1}},{{\theta}_{2}},\cdots,{{\theta}_{n}}\right\}$ and the associated density matrix is $\rho$. Then QB entropy of CBBA is defined as follows [43]:
${{E}_{Q}}(\mathbb{M})=S(\rho)+\sum\limits_{i\neq j}{Q{{I}_{ij}}},$ (25)
$S(\rho)=\sum\limits_{{{\theta}_{k}}\in\Theta}{\left|P\left({{\theta}_{k}}\right)\log_{2}P({{\theta}_{k}})\right|},$ (26)
where $P({{\theta}_{k}})=\operatorname{Tr}({{\mathbb{M}}_{{{\theta}_{k}}}}^{+}{{\mathbb{M}}_{{{\theta}_{k}}}}\rho)$.